Consider a sampling distribution with p=0.09 and samples of size n each. using the appropriate formulas, find the mean and the standard deviation of the sampling distribution of the sample proportion of the following parts:
A)for random sample of size n=5000
B)for random sample of size n=1000
C)for random sample of size n=500
Answer #1

Mean of sampling distribution od sample proportion = \mu\hat{p}

Standard deviation of sampling distribution of sample proportion = \sigma\hat{p}

A)

\mu\hat{p} = p = 0.09

\sigma\hat{p} = Sqrt( p ( 1 -p) / n)

= Sqrt(0.09 * 0.91 / 5000)

= 0.00405

B)

n= 1000

\mu\hat{p} = p = 0.09

\sigma\hat{p} = Sqrt( p ( 1 -p) / n)

= Sqrt(0.09 * 0.91 / 1000)

= 0.00905

C)

n = 500

\mu\hat{p} = p = 0.09

\sigma\hat{p} = Sqrt( p ( 1 -p) / n)

= Sqrt(0.09 * 0.91 / 500)

= 0.01280